# Measurement, and Mitigation

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of tasks, yet they also exhibit memorization of their training data. This phenomenon raises critical questions about model behavior, privacy risks, and the boundary between learning and memorization. Addressing these concerns, this paper synthesizes recent studies and investigates the landscape of memorization, the factors influencing it, and methods for its detection and mitigation. We explore key drivers, including training data duplication, training dynamics, and fine-tuning procedures that influence data memorization. In addition, we examine methodologies such as prefix-based extraction, membership inference, and adversarial prompting, assessing their effectiveness in detecting and measuring memorized content. Beyond technical analysis, we also explore the broader implications of memorization, including the legal and ethical implications. Finally, we discuss mitigation strategies, including data cleaning, differential privacy, and post-training unlearning, while highlighting open challenges in balancing the minimization of harmful memorization with utility. This paper provides a comprehensive overview of the current state of research on LLM memorization across technical, privacy, and performance dimensions, identifying critical directions for future work.
@@ 1 Introduction
Throughout the past few years, we have observed great strides in the capabilities of LLMs driven by changes in model architecture, training methodologies, and computational resources (Radford et al., 2018; Brown et al., 2020; Chowdhery et al., 2023; Naveed et al., 2023; Touvron et al., 2023; Wei et al., 2023). These advances have significantly improved natural language understanding, generation, and reasoning abilities, enabling these models to perform increasingly complex tasks across diverse domains (Alowais et al., 2023; Khurana et al., 2023; Minaee et al., 2025). However, in addition to their strong capabilities, LLMs continue to manifest critical limitations when evaluated on criteria such as privacy and security (Chang et al., 2024; Yao et al., 2024a; Das et al., 2025).
@@
With the advent of data scale used to train LLMs, the risks of privacy leakages through memorization have dramatically increased, as models trained on massive datasets containing potentially sensitive information may inadvertently reproduce verbatim passages from training data when prompted with specific triggers, compromising individual privacy and confidentiality without explicit consent (Dwivedi et al., 2023; Smith et al., 2023; Abdali et al., 2024). Thus, data memorization emerges as a critical vulnerability in LLMs, presenting significant privacy risks, including potential exposure of personally identifiable information (PII), copyright violations, and unauthorized reproduction of sensitive content (Carlini et al., 2021; Cremer et al., 2022; Clusmann et al., 2023; Karamolegkou et al., 2023; Song et al., 2023; Freeman et al., 2024; Hua et al., 2024; Morris et al., 2025).
This investigation explores memorization mechanisms in LLMs, examining contributing factors, detection methodologies, measurement approaches, and mitigation techniques. Unlike previous surveys on security and privacy (Siau and Wang, 2020; Fui-Hoon Nah et al., 2023; Guo et al., 2023; Hadi et al., 2023; Karabacak and Margetis, 2023; Min et al., 2023; Zhao et al., 2023; Zhu et al., 2023; Chang et al., 2024; Liu et al., 2024a; Myers et al., 2024; Raiaan et al., 2024; Stahl and Eke, 2024; Zhao et al., 2024), we present a focused examination of memorization phenomena in LLMs, dissecting the fundamental components contributing to and mitigating unintended data retention (Bender et al., 2021; Zhao et al., 2022a; Hartmann et al., 2023; Satvaty et al., 2024). This paper gives an in-depth review of the existing research on LLM memorization and its outcomes. Opposed to the survey by Satvaty et al. (2024), our analysis examines the technical underpinnings, evaluation methods, and privacy implications of memorization while providing concrete research directions. Figure 1 provides a visual taxonomy that maps and categorizes memorization, detection & mitigation of memorization, and corresponding risks.
@@ 2 Defining memorization
Agnostic to LLMs, memorization occurs when a model can reproduce specific training sequences when provided with appropriate contextual inputs (Carlini et al., 2023; Cooper et al., 2023; Schwarzschild et al., 2024). Currently, several definitions are used to quantify memorization in LLMs, including memorization defined by string matching (exact or approximate) or by loss differences.
2.1 Exact memorization
Verbatim memorization refers to a model’s exact reproduction of training data, often arising due to the high duplication of specific examples or overfitting (Carlini et al., 2021). This phenomenon occurs when an LLM, instead of generating novel responses or synthesizing information, simply retrieves and reproduces specific text snippets word-for-word from the data it was exposed to.
Perfect memorization describes a model that has exactly recorded its training data, assigning a probability only to inputs it has previously seen (Kandpal et al., 2022). When such a model generates an output, the process is mathematically equivalent to randomly selecting examples from its training dataset with a uniform probability.
Eidetic memorization (Carlini et al., 2021) occurs when a prompt $p$ causes the model to reproduce a verbatim string $s$ from training data. A stricter variant, $k$-eidetic memorization (Carlini et al., 2021), occurs when $s$ appears in no more than $k$ training examples, yet can still be extracted given the right prompt.
Discoverable memorization. A string $s$ is discoverably memorized (Carlini et al., 2023; Nasr et al., 2023) when, given a training example composed of prefix $p$ and suffix $s$, a LLM exactly reproduces the continuation $s$ that followed $p$ in the training data.
2.2 Approximate memorization
Approximate/Paraphrased memorization (Ippolito et al., 2023) is when LLMs generate outputs similar to training data in content, structure, or phrasing without exact replication. This differs from exact copying through variations, paraphrasing, or partial overlaps. The authors detect memorization by calculating edit distance between a generation and a target string, normalized by length, and choosing a threshold that determines when a sequence is deemed an approximate match.
2.3 Prompt-based memorization
Extractable memorization occurs when, without access to the training data, there exists a constructable prompt that invokes the model to generate an exact example from its training set (Nasr et al., 2023).
$k$-extractable memorization represents a stricter form of extractable memorization. A suffix is $k$-extractable (Biderman et al., 2023). when, given only its corresponding $k$-token prefix, the model reproduces the entire suffix verbatim. Unlike verbatim memorization, which manifests as direct replication, $k$-extractable memorization captures a model’s ability to retrieve and complete specific training sequences when prompted with only partial context.
($n,p$)-discoverable extraction (Hayes et al., 2025) formalizes the likelihood of retrieving a memorized string via repeated sampling. A string is ($n,p$)-discoverable if it appears in at least one of $n$ completions with probability $\geq p$.
2.4 Influence-based memorization
Counterfactual memorization (Feldman and Zhang, 2020; Zhang et al., 2023) quantifies how much a model’s predictions change based on whether a model is trained on a particular training example. For example, this can be instantiated as how much the loss changes on a datapoint when seen by a model during training versus not. Pappu et al. (2024) instantiate this definition by classifying datapoints as memorized when both a model trained on the datapoint achieves an approximate match under prefix-based decoding and a model not trained on that datapoint does not achieve an approximate match. As it is computationally prohibitive to retrain a model per datapoint-exlusion, Zhang et al. (2023) and Pappu et al. (2024) train a smaller number of models that exclude entire subsets of data used to test for memorization.
@@ 3 Factors influencing memorization
With the increasing scale and capability of LLMs, a range of training and inference-time factors have been identified as influencing memorization behavior. This section systematically analyzes the primary variables that govern memorization.
@@
Model parameter size. Model size strongly correlates with increased memorization in LLMs, with larger models demonstrating both greater capacity to retain training data and increased vulnerability to extraction attacks. Carlini et al. (2021) first introduced the relationship between model size and memorization, observing that memorization scales log-linearly with model size. Subsequent research has consistently reinforced this finding across multiple dimensions of memorization. For instance, Tirumala et al. (2022) observed that larger LLMs not only memorize more content but do so more rapidly during training, a phenomenon that cannot be fully explained by conventional overfitting or hyper-parameter tuning. This enhanced memorization capability directly impacts privacy concerns, as Li et al. (2024) demonstrated that extraction attacks become significantly more effective against larger LLMs due to their advanced memorization capabilities. The scaling effect has been further corroborated by (Lee et al., 2022; Tirumala et al., 2022; Carlini et al., 2023; Freeman et al., 2024; Kiyomaru et al., 2024; Li et al., 2024; Hayes et al., 2025; Morris et al., 2025).
@@
Training data duplication. Duplicate examples in training data skew LLMs toward overrepresented content, diminishing output diversity and increasing verbatim reproduction of training material (Carlini et al., 2021). As demonstrated in Lee et al. (2022), de-duplication substantially reduces memorization, with models trained on original data exhibiting a tenfold increase in memorized token generation compared to those trained on datasets, where exact substrings and near-duplicates were removed. Exploring further, Kandpal et al. (2022) identified a superlinear relationship between training data duplication and memorization in LLMs, where rarely duplicated training samples are seldom memorized. However, models generate duplicated sequences at rates significantly lower than perfect memorization would predict.
@@
Sequence length. Longer sequences increase memorization in LLMs. Carlini et al. (2023) found that memorization increases logarithmically with sequence length, with the verbatim reproduction probability rising by orders of magnitude as sequences extend from 50 to 950 tokens. Similarly, extraction methodologies, including continuous soft prompting and dynamic soft prompting techniques, demonstrate a consistent pattern wherein the volume of extracted data grows proportionally with increasing prefix token sizes (Wang et al., 2024).
@@
Tokenization. Models trained with larger Byte Pair Encoding (BPE) vocabularies were found to memorize significantly more sequences from their training data. Kharitonov et al. (2021) studied how different vocabulary sizes, produced via BPE, affect a transformer LLMs’ tendency to memorize sequences. Memorization was particularly strong for named entities, URLs, and uncommon phrases, often becoming single tokens under larger BPE settings.
@@
Sampling Methods. Sampling methods significantly impact the extraction of memorized content from LLMs, with stochastic approaches consistently outperforming greedy decoding in revealing memorized training data. While Carlini et al. (2021) initially established that greedy decoding can reveal memorized sequences but misses others due to its low diversity, subsequent research has demonstrated the superior effectiveness of other sampling methods. Yu et al. (2023) showed that optimizing sampling parameters like top-k, nucleus sampling, and temperature can substantially increase memorized data extraction, in some cases doubling previous baselines. This finding was reinforced by Tiwari and Suh (2025), who discovered that randomized decoding nearly doubles leakage risk compared to greedy decoding, contradicting earlier assumptions. Hayes et al. (2025) provided a theoretical foundation for these observations by introducing a probabilistic framework that demonstrates how repeated sampling with varied decoding parameters can expose memorization hidden under greedy approaches. Given that no single decoding method minimizes leakage across all scenarios, Tiwari and Suh (2025) emphasized the importance of assessing memorization under diverse sampling strategies.
@@
While data duplication has been broadly associated with increased memorization, it remains uncertain whether this effect systematically varies across different data modalities (e.g., code, prose, etc.). Additionally, although inference hyper-parameters, such as decoding methods, are known to impact memorization, existing techniques do not reliably prevent the extraction of memorized content. This underscores the need for decoding strategies explicitly designed to minimize memorization. In contrast, the influence of training hyper-parameters on memorization remains poorly understood, particularly with regards to the role of overfitting through learning rate and regularization mechanisms. Moreover, the temporal dynamics of memorization in the presence of data duplication remain underexplored. Specifically, it is unclear how the timing of when duplicated examples are encountered during training affects memorization likelihood.
Open Questions 1. How do data duplication and data type affect memorization extent? 2. What is required to design a decoding method that minimizes leakage? 3. What factors cause larger LLMs to memorize more? 4. How can we distinguish useful generalization from memorization in LLMs? 5. How do training hyperparameters influence memorization? 6. Does the temporal order of data presentation during training affect memorization likelihood?

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of tasks, yet they also exhibit memorization of their training data. This phenomenon raises critical questions about model behavior, privacy risks, and the boundary between learning and memorization. Addressing these concerns, this paper synthesizes recent studies and investigates the landscape of memorization, the factors influencing it, and methods for its detection and mitigation. We explore key drivers, including training data duplication, training dynamics, and fine-tuning procedures that influence data memorization. In addition, we examine methodologies such as prefix-based extraction, membership inference, and adversarial prompting, assessing their effectiveness in detecting and measuring memorized content. Beyond technical analysis, we also explore the broader implications of memorization, including the legal and ethical implications. Finally, we discuss mitigation strategies, including data cleaning, differential privacy, and post-training unlearning, while highlighting open challenges in balancing the need to minimize harmful memorization with model utility. This paper provides a comprehensive overview of the current state of research on LLM memorization across technical, privacy, and performance dimensions, identifying critical directions for future work.
@@ 1 Introduction
Throughout the past few years, we have observed great strides in the capabilities of LLMs driven by changes in model architecture, training methodologies, and computational resources [radford2018improving, brown2020languagemodelsfewshotlearners, chowdhery2022palmscalinglanguagemodeling, naveed2023comprehensive, touvron2023llamaopenefficientfoundation, wei2023chainofthoughtpromptingelicitsreasoning]. These advances have significantly improved natural language understanding, generation, and reasoning abilities, enabling these models to perform increasingly complex tasks across diverse domains [alowais2023revolutionizing, khurana2023natural, minaee2025largelanguagemodelssurvey]. However, in addition to their strong capabilities, LLMs continue to manifest critical limitations when evaluated on criteria such as privacy and security [10.1145/3641289, yao2024survey, das2025security].
@@
With the advent of data scale used to train LLMs, the risks of privacy leakages through memorization have dramatically increased, as models trained on massive datasets containing potentially sensitive information may inadvertently reproduce verbatim passages from training data when prompted with specific triggers, compromising individual privacy and confidentiality without explicit consent [dwivedi2023opinion, smith2023identifying, abdali2024securing, nasr2023scalableextractiontrainingdata]. Thus, data memorization emerges as a critical vulnerability in LLMs, presenting significant privacy risks, including potential exposure of Personally Identifiable Information (PII), copyright violations, and unauthorized reproduction of sensitive content [carlini2021extractingtrainingdatalarge, cremer2022cyber, clusmann2023future, karamolegkou2023copyright, song2023digital, freeman2024exploringmemorizationcopyrightviolation, hua2024limitations, morris2025much].
This investigation explores memorization mechanisms in LLMs, examining contributing factors, detection methodologies, measurement approaches, and mitigation techniques. Unlike previous surveys on security and privacy [siau2020artificial, fui2023generative, guo2023evaluating, hadi2023large, karabacak2023embracing, min2023recent, zhao2023survey, zhu2023large, 10.1145/3641289, liu2024datasets, myers2024foundation, raiaan2024review, stahl2024ethics, zhao2024explainability], we present a focused examination of memorization phenomena in LLMs, dissecting the fundamental components contributing to and mitigating unintended data retention [bender2021dangers, zhao2022provably, hartmann2023sok, satvaty2024undesirable]. Our work provides unique contributions that distinguish it from existing surveys. We significantly expand coverage of existing work presented in [hartmann2023sok, satvaty2024undesirable]. While Hartmann, et al. [hartmann2023sok] focused on definitions, implications (e.g., model performance and alignment, privacy, security and confidentiality, copyright, and auditing), detection, and prevention, this paper provides an in-depth review of the existing research on LLM memorization and its outcomes by emphasizing concept differences and relationships, main research challenges with key barriers and research directions at different memorization stages, the key principles, requirements and practicality, and core limitations and statistical soundness of detection techniques, and research challenges for detection, mitigation, and privacy and legal risk of memorization. In contrast with the survey by [satvaty2024undesirable], our analysis examines the technical underpinnings, evaluation methods, and privacy implications of memorization and provides concrete research directions. We organize open research questions systematically by memorization subtopics and provide some rationale before presenting each question. While identifying open problems is inherently subjective, we select the most pressing issues based on current research challenges, though we acknowledge that other researchers may prioritize different questions and that these priorities will likely evolve as the field advances.
Figure 1 provides a visual taxonomy that maps and categorizes memorization, detection and mitigation of memorization, and corresponding privacy and legal risks.
@@ 2 Defining memorization
This research examines cases of unintended memorization, where LLMs reproduce sensitive, private, or verbatim training data. While memorizing facts or grammatical rules is essential for model utility, the inadvertent reproduction of private or non-public data raises significant privacy and ethical concerns. The central challenge, which this paper aims to systematize, is distinguishing between beneficial and harmful forms of memorization. This work focuses specifically on the latter, providing a comprehensive overview of the mechanisms, detection methods, and mitigation strategies for unintended data reproduction in LLMs.
@@
Agnostic to LLMs, memorization occurs when a model can reproduce specific training sequences when provided with appropriate contextual inputs [carlini2023quantifyingmemorizationneurallanguage, cooper2023report, schwarzschild2024rethinking]. This is the type of memorization most commonly studied in the context of generative models. Currently, several definitions are used to quantify memorization in LLMs, including memorization defined by string matching (exact or approximate) or by loss differences. In addition, Hayes, et al. [hayes2025measuringmemorizationlanguagemodels] employs baseline comparisons to distinguish genuine memorization from coincidental sequence generation.
However, it is important to recognize that memorization is not inherently negative; LLMs also depend on the beneficial memorization of factual, linguistic, and semantic regularities that underpin generalization, coherence, and factual accuracy. Specifically, memorization contributes positively to a model’s ability to store and retrieve knowledge. The critical distinction lies in what is memorized and how it is used.
Appendix’s Section 1 expands on these memorization definitions. The heterogeneity of these definitions across the literature has resulted in a fragmented landscape of detection and mitigation strategies. This lack of a unified evaluation framework is a key challenge, as existing definitions are constrained and exhibit considerable overlap, hindering the direct comparison of results across studies. To provide clarity, Table I systematizes these concepts, outlining their key differences, interrelationships, and contributions and effects.
Category Definition Defining Characteristic Relationship to Other Definitions Contribution & Effect
[…]
Effect: Forms the legal and privacy basis for concerns about copyright infringement and PII exposure.
[…]
Effect: Highlights the limitations of simple string-matching detection methods.
[…]
Effect: Signals a deeper, more concerning level of data retention by the model.
[…]
Effect: Models the worst-case potential for data extraction.
[…]
Effect: Enables systematic, large-scale privacy auditing and unwanted “auto-completion” risk analysis.
[…]
Effect: Used to analyze how prompt length affects memorization recall.
[…]
Effect: Better quantifies data exposure risk under diverse sampling strategies.
[…]
Effect: Allows researchers to causally separate memorization from learned generalization, which is crucial for unlearning methods and for legally defensible audits of data influence.
[…]
Effect: Potentially relevant for fair use legal arguments, by linking the “effort” (prompt complexity) to the “recall” (output length).
TABLE I: A Taxonomy of Memorization Definitions in Large Language Models
@@ 3 Factors influencing memorization
The memorization behavior of LLMs is governed by a range of factors spanning training and inference.
@@
Model parameter size. Model size strongly correlates with increased memorization in LLMs, with larger models demonstrating both greater capacity to retain training data and increased vulnerability to extraction attacks. Carlini, et al. [carlini2021extractingtrainingdatalarge] first introduced the relationship between model size and memorization, observing that memorization scales log-linearly with model size. Subsequent research has consistently reinforced this finding across multiple dimensions of memorization. For instance, Tirumala, et al. [tirumala2022memorizationoverfittinganalyzingtraining] observed that larger LLMs not only memorize more content but do so more rapidly during training, a phenomenon that cannot be fully explained by conventional overfitting or hyper-parameter tuning. This enhanced memorization capability directly impacts privacy concerns, as Li, et al. [li2024llmpbeassessingdataprivacy] demonstrated that extraction attacks become significantly more effective against larger LLMs due to their advanced memorization capabilities. The scaling effect has been further corroborated by [lee2022deduplicatingtrainingdatamakes, tirumala2022memorizationoverfittinganalyzingtraining, carlini2023quantifyingmemorizationneurallanguage, freeman2024exploringmemorizationcopyrightviolation, kiyomaru-etal-2024-comprehensive-analysis, li2024llmpbeassessingdataprivacy, hayes2025measuringmemorizationlanguagemodels, morris2025much].
@@
Training data duplication. Duplicate examples in training data skew LLMs toward overrepresented content, diminishing output diversity and increasing verbatim reproduction of training material [carlini2021extractingtrainingdatalarge]. As demonstrated in [lee2022deduplicatingtrainingdatamakes], de-duplication substantially reduces memorization; models trained on original data exhibited a tenfold decrease in memorized token generation compared to those trained on datasets where exact substrings and near-duplicates were not removed. Exploring further, Kandpal, et al. [kandpal2022deduplicatingtrainingdatamitigates] identified a superlinear relationship between training data duplication and memorization in LLMs, where rarely duplicated training samples are seldom memorized.
Current deduplication techniques encode a compromise between human intuition about content redundancy and the computational efficiency required at a massive scale. Consequently, studies show that memorization often persists after syntactic deduplication [lee2022deduplicatingtrainingdatamakes], driven primarily by near-duplicates (e.g., paraphrased sentences or text with minor edits) that fall below the similarity thresholds currently in use [SHSP2024]. This suggests the need for training data attribution–based approaches that can analyze the influence of near-duplicates and identify which content segments the model internally treats as equivalent, rather than relying solely on surface-level hash or substring matching. Adopting such a model-centric perspective would redefine duplication based on the learned internal representations of the model itself, rather than on pre-defined similarity heuristics. For example, recent studies have successfully used influence functions [KohLiang2017Understanding] to quantify the contribution of individual training instances (e.g., near-duplicates) to specific memorized outputs, confirming that models can treat syntactically distinct but semantically related data as equivalent [menta2025analyzing]. Understanding this internal behavior is crucial, as it helps explain why even highly duplicated sequences are generated at rates significantly lower than perfect memorization would predict, suggesting that other complex dynamics, not just data frequency, govern the recall process.
@@
Sequence length. Longer sequences increase memorization in LLMs. Carlini, et al. [carlini2023quantifyingmemorizationneurallanguage] found that memorization increases logarithmically with sequence length, with the verbatim reproduction probability rising by orders of magnitude as sequences extend from 50 to 950 tokens. Specifically, Carlini, et al. [carlini2023quantifyingmemorizationneurallanguage] showed that the length of the prefix used as a prompt matters. Similarly, extraction methodologies, including continuous soft prompting and dynamic soft prompting techniques, demonstrate a consistent pattern wherein the volume of extracted data grows proportionally with increasing prefix token sizes [wang-etal-2024-unlocking].
@@
Tokenization. Models trained with larger Byte Pair Encoding (BPE) vocabularies were found to memorize significantly more sequences from their training data. [kharitonov2021bpe] studied how different vocabulary sizes, produced via BPE, affect a transformer LLMs’ tendency to memorize sequences. Memorization was particularly strong for named entities, URLs, and uncommon phrases, often becoming single tokens under larger BPE settings.
@@
Sampling methods. Sampling methods significantly impact the extraction of memorized content from LLMs, with stochastic approaches consistently outperforming greedy decoding in revealing memorized training data. While Carlini, et al. [carlini2021extractingtrainingdatalarge] initially established that greedy decoding can reveal memorized sequences but misses others due to its low diversity, subsequent research has demonstrated the superior effectiveness of other sampling methods. Yu, et al. [yu2023bag] showed that optimizing sampling parameters like top-k, nucleus sampling, and temperature can substantially increase memorized data extraction, in some cases doubling previous baselines. This finding was reinforced by Tiwari and Suh [tiwari2025sequencelevelleakagerisktraining], who discovered that randomized decoding nearly doubles leakage risk compared to greedy decoding, contradicting earlier assumptions. Hayes, et al. [hayes2025measuringmemorizationlanguagemodels] provided a theoretical foundation for these observations by introducing a probabilistic framework that demonstrates how repeated sampling with varied decoding parameters can expose memorization hidden under greedy approaches. Given that no single decoding method minimizes leakage across all scenarios, Tiwari and Suh [tiwari2025sequencelevelleakagerisktraining] emphasized the importance of assessing memorization under diverse sampling strategies.
Data modality and content characteristics. The influence of data duplication and other factors is not uniform across different types of data. Research into multi-modal systems shows that memorization risks vary across modalities like text, images, and audio [aghabagherloo2025impact]. For example, diffusion models have been shown to replicate images from their training set, a form of visual memorization [somepalli2023understanding], and vision-language models can exhibit cross-modality memorization, where a textual prompt elicits a memorized image or vice-versa [wen2025quantifyingcrossmodalitymemorizationvisionlanguage]. Beyond modality, the inherent characteristics of the data play a crucial role. Unique or novel sequences, such as PII or secret keys, are more susceptible to memorization because the model cannot generalize from other examples and must store the information explicitly to minimize loss on that data point[feldman2020neural].
Interplay of factors. The factors influencing memorization are intertwined, creating a complex system where their effects are often synergistic and non-stochastic. This process can be imagined as a cascade with data duplication providing a strong initial signal, which is then captured by the large parameter size of models, a relationship well-described by scaling laws [hernandez2022scaling]. This is exacerbated by tokenization, which can collapse lengthy, unique sequences into single, easily memorized units, directly impacting how sequence length contributes to memorization risk. Specific prompts act as keys while varied sampling methods can either reveal the most obvious memorized content or, over repeated trials, uncover a far wider range of subtly store information [hayes2025measuringmemorizationlanguagemodels]. As further discussed in Section 4, these dynamics can be unified through the lens of data compression, where models learn to memorize predictable sequences as an efficient strategy to minimize training loss [schwarzschild2024rethinking]. This latent memorization is shaped by the temporal training dynamics that favor recently seen data [jagielski2023measuring]. This intricate interplay of factors means that memorization often arises from a “perfect storm” of factors, underscoring that various confounding factors may determine the strength of memorization as well as that effective mitigation requires a holistic approach rather than addressing any single factor in isolation.
@@
While data duplication has been broadly associated with increased memorization, it remains uncertain whether this effect systematically varies across different data modalities (e.g., code, prose). Additionally, although inference hyper-parameters, such as decoding methods, are known to influence memorization, existing techniques do not reliably prevent the extraction of memorized content. This underscores the need for decoding strategies explicitly designed to minimize memorization. In contrast, the influence of training hyper-parameters on memorization remains poorly understood, particularly with regards to the role of overfitting through learning rate and regularization mechanisms. Pappu, et al. [pappu2024measuringmemorizationrlhfcode] noted that larger learning rates can increase memorization, but within the narrow domain of Reinforcement Learning from Human Feedback (RLHF) for code they investigate. Whether this is simply equivalent to whether the model well-fits the training data (i.e., under or overfits) is unknown. Moreover, the temporal dynamics of memorization in the presence of data duplication remain underexplored. Specifically, it is unclear how the timing of when duplicated examples are encountered during training affects memorization likelihood.
Open Questions 1. Beyond duplication, what intrinsic data properties govern memorization? 2. What is the optimal trade-off between memorization prevention vs. utility at inference? 3. How does model architecture scale influence the mechanisms of memorization? 4. How can we distinguish useful generalization from memorization in LLMs?
OQ1. Since models memorize not only duplicated but also unique sequences, memorization risk appears linked to intrinsic data properties like low complexity or high predictability. The core research challenge is to formalize a predictive theory of this “memorization potential,” enabling proactive data curation to identify and remove high-risk content prior to training.
OQ2. Decoding-time defenses rely on a real-time “leakage score” to guide text generation away from memorized training data [ippolito-etal-2023-preventing]. The open frontier is to formulate a metric that is both computationally tractable for inference and can navigate the fundamental trade-off between preventing harmful regurgitation and maintaining model utility.
OQ3. It remains unclear if larger models simply have more capacity for memorization or if the internal mechanisms themselves evolve with scale. Tracing “memorization circuits” [meng2022locating] within components like FFN layers theorized to act as key-value memories [geva2022transformer] is crucial for understanding how these circuits change in smaller models and for designing future architectures that are inherently safer.
OQ4. The ideal test for memorization involves a counterfactual [zhang2023counterfactual]: would the model still produce this output if this exact training example were removed? Since retraining a massive model to answer this is computationally impossible, the critical research frontier is developing scalable approximations that provide a practical, operational line between useful knowledge and harmful regurgitation.